Given the utility function U =10Q1Q2 and relative prices of P1= ₦20, P2 = ₦10 and consumer money income of ₦400. You are required to determine the units of each commodity to consume at the equilibrium point.

   


1
Expert's answer
2022-02-25T10:46:58-0500

"U= 10Q_1Q_2"

"P_1= 20, P_2= 10, I= 400"

Budget line= "20Q_1+10Q_2= 400"

At equilibrium,

"\frac{Mu_1}{MU_2}= \frac{P_1}{P_2}"

"Mu_1= \frac{\delta U}{\delta Q_1}= 10Q_2"

"Mu_2= \frac{\delta U}{\delta Q_2}= 10Q_1"


"\frac{10Q_2}{10Q_1}= \frac{20}{10}"

"Q_1= 0.5Q_2"

"Q_2= 2Q_1"

Plug the these values in the budget Equation

"20(0.5Q_2)+10Q_2= 400"

"20Q_2= 400"

"Q_2^*= 20"


"20Q_1+10(2Q_1)= 400"

"40 Q_1= 400"

"Q_1^*= 10"